Assume that the total score (from both teams) for college football games averages μ = 42 points per
game, and that the distribution of total points is approximately normal with σ = 20.

b. What proportion of college football games have a point total between 20 and 60?
1
Expert's answer
2020-05-31T18:27:02-0400

Let "X" has normal distribution "X"~"N(42,20^2)", then "Z=\frac{X-42}{20}"~"N(0,1)",

"P(20<X<60)=\\
P(\frac{20-42}{20}<\frac{X-42}{20}<\frac{60-42}{20})=\\
P(-1.1<Z<0.9)=\\
P(Z<0.9)-P(Z\leq -1.1)\approx\\
0.8159-0.1357=0.6802."

Approximately "100*0.6802"%"=68.02"% of games have a point total between 20 and 60.

Answer: 68.02%.